# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

### v2

@@ Abstract
We present ACDiT, a novel Autoregressive blockwise Conditional Diffusion Transformer, that innovatively combines autoregressive and diffusion paradigms for modeling continuous visual information. By introducing a block-wise autoregressive unit, ACDiT offers a flexible interpolation between token-wise autoregression and full-sequence diffusion, bypassing the limitations of discrete tokenization. The generation of each block is formulated as a conditional diffusion process, conditioned on prior blocks. ACDiT is easy to implement, as simple as creating a Skip-Causal Attention Mask (SCAM) on standard diffusion transformer during training. During inference, the process iterates between diffusion denoising and autoregressive decoding that can make full use of KV-Cache. We show that ACDiT performs best among all autoregressive baselines under similar model scales on image and video generation tasks. We also demonstrate that benefiting from autoregressive modeling, pretrained ACDiT can be transferred in visual understanding tasks despite being trained with the diffusion objective. The analysis of the trade-off between autoregressive modeling and diffusion demonstrates the potential of ACDiT to be used in long-horizon visual generation tasks. We hope that ACDiT offers a novel perspective on visual autoregressive generation and unlocks new avenues for unified models.
@@ 1 Introduction
The concept of predicting the future has been a fundamental principle of Artificial Intelligence, ranging from the “next token prediction” objective of autoregressive language modeling [55, 6, 1] to predicting the next action in Reinforcement Learning [48, 62, 12, 57]. The success of Large Language Models (LLMs) [70, 71, 6, 1, 34, 2, 80] exemplify the power of such philosophy, demonstrating that complex abilities can arise from this simple objective.
@@
However, the multimodal aspects of information beyond language have yet to fully capitalize on this paradigm. In the realm of visual generation, diffusion models [31, 63, 14, 64] have demonstrated superior generative capabilities, producing creative outputs that are virtually indistinguishable from human-generated content, as evidenced by innovations like Sora [5] and Stable Diffusion [58, 53, 20]. Notwithstanding their remarkable achievements, these models operate in a non-autoregressive manner. Diffusion models receive corrupted target sequences with complete length and reconstruct the intended output through in-place iterative refinement. The term “in-place” underscores the distinct nature of this approach compared to the autoregressive model, which provides subsequent prediction, thereby extending the sequence and progressing towards the future. Such variation makes the model struggle to learn temporal correlation in vision data and poses difficulties for developing integrated frameworks capable of seamlessly bridging the vision foundation model with unified multimodality modeling [17, 68, 87, 75, 79, 78] and world model [81, 19, 7, 88, 36, 77].
Existing works attempt to build the visual autoregressive model by converting visual generation tasks into discrete token prediction tasks with vector quantization techniques [21, 68, 75, 69] and training with a next-token prediction objective. However, approaching the continuous distribution requires huge vocabulary sizes and a high utilization rate [85, 76], which is a complex objective.
@@
ACDiT is easy to implement, as simple as adding a Skip-Causal Attention Mask to the current diffusion transformer [52] architecture during training. The inference process is formatted as an iteration between the conditional diffusion denoising process within a block, conditioned on the complete clean context, and autoregressive generation of a new block appended as the new context. In this way, KV-Cache can be used for faster inference. In general, ACDiT offers the following inherent advantages: (i) ACDiT simultaneously learns the causal interdependence across blocks with autoregressive modeling and the non-causal dependence within blocks with diffusion modeling. (ii) ACDiT is endowed with clean continuous visual input, which eliminates the requirement for vector quantization and improves the transfering to visual understanding tasks after being trained with generative objectives. (iii) ACDiT makes full use of KV-Cache for flexible autoregressive generation in any length and can potentially benefit from other latest long-context techniques in text for long video generation.
We validate ACDiT in both image and video generation tasks on ImageNet [60] and UCF101 [65], respectively. Experimental results demonstrate that ACDiT outperforms all autoregressive baselines of the same model scale and achieves visual quality comparable to full-sequence diffusion models while exhibiting significantly higher inference speed when extended to long sequences. We hope ACDiT presents a fresh perspective on visual autoregressive generation and paves the way for exploring unified multimodal and world models. The codes and models are available at https://github.com/thunlp/ACDiT.
@@ 2 Related Work
Diffusion Models. The field of image generation has witnessed remarkable advancements with the introduction of diffusion models [31, 63, 14, 49]. U-Net [59] is the early mainstream choice of network architecture [64, 49, 58, 53]. Following that, Transformer [73] is applied to diffusion models for image generation, with groundbreaking work such as DiT [52] and U-ViT [3] marking significant milestones. A series work, including PixArt-$\alpha,\delta,\Sigma$[9, 11, 10], demonstrate the capability of DiT on text-to-image tasks. Additionally, several studies have applied DiT to video generation, such as Lumiere [4] and Movie Gen [54].
@@
Autoregressive Generation on Discrete Tokens. The iGPT [13] first proposes autoregressively generating raw image pixels as a raster-scan sequence. Most visual autoregressive models [56, 75, 83, 15, 16] follow VQGAN [21], which pioneers this direction by training an autoregressive transformer on discrete tokens produced by VQVAE [72]. LlamaGen [67] enhances the image tokenizer and scales up the autoregressive transformers building on the latest Llama architecture [70]. Inspired by RQ-Transformer [38], VAR [69] proposes the next-scale prediction and obtain good improvement. Some subsequent works [69, 43, 82] extend next-scale prediction to text-to-image generation.
@@
Autoregressive Generation on Continous Features. Some recent works make some early exploration on achieving visual autoregressive generation without discrete tokens and combining the advantage of diffusion and autoregressive. Diffusion Forcing [8] trains a causal autoregressive model to generate blocks without fully diffusing past ones and implement it on small RNN. MAR [41] proposes the diffusion loss to learn the autoregressive conditional distribution on the head of the main Transformer with a small MLP network. Transfusion [87] and Monoformer [86] conduct joint training with language modeling loss and diffusion loss in a single transformer, while within image generation, their design is identical to the standard diffusion model. Dart [25] proposes to autoregressively denoised the complete image within in a Transformer. Different from these work, ACDiT redefine the autoregressive unit and generates each block based on clear past, providing a flexible framework for both image and video generation.
@@
Figure 2: (a): for each noised block $n_{i}$, it can only attend previous clean latent $c_{0},c_{1},\dots,c_{i-1}$ and itself. Each clean block $c_{i}$ only attends the previous clean latent block. (b): ACDiT can effectively utilize the KV-Cache for autoregressive inference. (c): the 3D view of ACDiT, where $B,L,T$ denotes the block size, number of blocks, and denoising time steps, respectively. The darker color denotes higher noises.
3 Prerequiste
3.1 Autoregressive Modeling
Autoregression asserts that the value at each timestep is contingent upon its preceding values. This principle is exemplified in autoregressive language models, which iteratively predict the probability distribution of subsequent tokens. Given a sequence of tokens ($x_{1},x_{2},\dots,x_{n}$), a salient characteristic of autoregression is that the prediction of $x_{i}$ is only dependent on its prefix ($x_{1},x_{2},\dots,x_{i-1}$). Upon determining $x_{i}$, it is concatenated with the preceding sequence, thereby forming the conditioning context ($x_{1},x_{2},\dots,x_{i}$) for predicting $x_{i+1}$. Therefore, the likelihood of sequence can be factorized as:
$p(x_{1},x_{2},\dots,x_{n})=\prod_{i=1}^{n}p(x_{i+1}|x_{1},x_{2},\dots,x_{i}).$ (1)
Thanks to the flexibility of self-attention in Transformers, autoregressive models can be effectively implemented by adding a causal mask in Transformer attention block [73].
3.2 Diffusion
Diffusion models, in contrast, conceptualize a noise-infusion and denoising process, which is defined by gradually adding noise to the initial data $n_{0}$ and training the model to learn the inverse mapping. Formally, the noised data $n^{(t)}$ at each step $t$ is sampled by $q(n^{(t)}|n^{(t-1)})=\mathcal{N}(x_{t};\sqrt{\alpha^{(t)}}n^{(0)},(1-\alpha^{(t)})\mathbf{I})$ [^1] For clarification, we use subscript t to denote timesteps in autoregressive models and superscript (t) to denote the timesteps in diffusion models., which is equivalent to add a Guassain noise to the previous samples: $n^{(t)}=\sqrt{\alpha^{(t)}}n^{(t-1)}+\sqrt{1-\alpha^{(t)}}\epsilon^{(t)},\epsilon^{(t)}\sim\mathcal{N}(0,\mathbf{I})$ while the diffusion model $p_{\theta}$ is trained to learn the reverse process $p_{\theta}(n^{(t-1)}|n^{(t)})=\mathcal{N}(\mu_{\theta}(n^{(t)}),\beta^{(t)}\mathbf{I})$. With the reparameterization trick, the network $\mu_{\theta}(n^{(t)})$ can be reparameterized as noise prediction network $\epsilon_{\theta}(n^{(t)})$ and the training objective can be simple as:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\sim U[0,1],\epsilon\sim\mathcal{N}(0,I)}||\epsilon_{\theta}(n^{(t)},t)-\epsilon^{(t)}||_{2}.$ (2)
During the inference phase, the denoising process is initialized with a random Gaussian noise sample $n^{(T)}$, followed by $T^{\prime}$ denoising steps, ultimately yielding a single deterministic samples $\tilde{n}^{(0)}$ from its underlying distribution. Typically, the denoising process in diffusion models operates “in-place”, meaning that each new denoising step directly replaces the previous step’s input. This differs from autoregressive modeling, where the value of a subsequent step is appended to the existing sequence.
3.3 Desiderata
A robust autoregressive diffusion method should integrate the strengths of both autoregressive modeling and diffusion. To achieve this synergy, we have identified three critical desiderata that the framework must meet:
  1. 1.
  2. 2.
    Both the autoregressive modeling and the denoising process should optimally utilize the entire parameter space of the neural network. In an elegant fusion of autoregressive models and diffusion, neither component should be relegated to an auxiliary role. Instead, they should function as integral, complementary elements of the system.
  3. 3.
    The denoising process should directly attend comprehensively to the entire sequence of past sequences. Failure to do so would necessitate that the denoise process’s condition encapsulates all preceding information, placing an unrealistic demand on the lossless compression capabilities of the feature space. On the contrary, a holistic input of all past information in each denoise step ensures a more effective processing of temporal dependencies.
Based on these desiderata, we analyze representative existing autoregressive diffusion methods. Diffusion-Forcing [8] proposes to use different-level random noise in the different positions of a sequence. Thus, in the inference process, denoising the subsequent positions from a clean past can be seen as a special case denoising different levels of noise. Their method does not meet the first desideratum. In the training process, the future is not predicted from the precise representation of the past. In MAR [41], the diffusion process is trained based on the latent in the last position, which does not satisfy the third desideratum. Moreover, the diffusion process sorely leverages the head part of the network, which conflicts with the second desideratum. In Transfusion [87] and Monoformer [86], despite that the diffusion utilizes the full network parameters for both autoregressive modeling and diffusion, it does not utilize the “predicting the future” objective within the visual information. Moreover, when trained with multiple visual input, such as multiple images, the latter image will attend to the noise version of the former image. To comprise this deficiency, they half the noise schedule, i.e., limiting maximum noise steps to 500 from 1000, in 20% image captioning pairs, which is not an optimal strategy.
Figure 3: Left: an illustrative figure of the ideal dependencies between the noise and clean information, which is adopted by ACDiT. Right: the dependencies of diffusion head in MAR.
4 ACDiT

### v3

@@ Abstract
Autoregressive and diffusion models have achieved remarkable progress in language models and visual generation, respectively. We present ACDiT, a novel Autoregressive blockwise Conditional Diffusion Transformer, that innovatively combines autoregressive and diffusion paradigms. By introducing a block-wise autoregressive unit, ACDiT offers a flexible interpolation between token-wise autoregression and full-sequence diffusion, bypassing the limitations of discrete tokenization. The generation of each block is formulated as a conditional diffusion process, conditioned on prior blocks. ACDiT is easy to implement, as simple as applying a specially designed Skip-Causal Attention Mask on the standard diffusion transformer during training. During inference, the process iterates between diffusion denoising and autoregressive decoding that can make full use of KV-Cache. We validate the effectiveness of ACDiT on image, video, and text generation and show that ACDiT performs best among all autoregressive baselines on similar model scales on visual generation tasks. We also demonstrate that benefiting from autoregressive modeling, pretrained ACDiT can be transferred to visual understanding tasks despite being trained with the generative objective. The analysis of the trade-off between autoregressive and diffusion demonstrates the potential of ACDiT to be used in long-horizon visual generation tasks. We hope that ACDiT offers a novel perspective on visual autoregressive generation and sheds light on new avenues for unified models.
@@ 1 Introduction
Autoregressive modeling has been a central paradigm in artificial intelligence, most notably driving the success of large language models through next-token prediction (Touvron et al., 2023a; b; Brown, 2020; Achiam et al., 2023; Jiang et al., 2023; Bai et al., 2023; Yang et al., 2024). By decomposing complex distributions into a sequence of conditional predictions, autoregressive models provide a natural mechanism for modeling long-range dependencies and structured generation. Similar predictive formulations also underpin reinforcement learning (Mnih, 2013; Schulman et al., 2017; Chen et al., 2021; Reed et al., 2022) and world models (Ha and Schmidhuber, 2018), where future states are predicted conditioned on past observations. These successes motivate us to revisit autoregressive factorization beyond language and ask how its advantages can be effectively extended to high-dimensional visual generation.
@@
In the realm of visual generation, diffusion models (Ho et al., 2020; Song et al., 2020a; Dhariwal and Nichol, 2021; Song et al., 2020b) have demonstrated superior generative capabilities, producing creative outputs that are virtually indistinguishable from human-generated content, as evidenced by innovations like Sora (Brooks et al., 2024) and Stable Diffusion (Rombach et al., 2022; Podell et al., 2023; Esser et al., ). Despite their remarkable success, diffusion models operate in a non-autoregressive manner. They take as input corrupted target sequences of the full length and reconstruct the intended output through in-place iterative refinement. The term “in-place” highlights a fundamental distinction from autoregressive models, which provide subsequent prediction, thereby extending the sequence and progressing towards the future. Such variation makes the diffusion model less effective at capturing long-range temporal dependencies in visual data and poses difficulties for developing integrated frameworks capable of seamlessly bridging the vision foundation model with unified multimodality modeling (Dong et al., 2024; Team, 2024; Zhou et al., 2024a; Wang et al., 2024b; Xie et al., 2024; Wu et al., 2024b) and world model (Yang et al., 2023; Du et al., 2023; Bruce et al., 2024; Zhou et al., 2024b; Wu et al., 2024a).
To build the visual autoregressive model, existing works convert visual generation tasks into discrete token prediction tasks with vector quantization techniques (Esser et al., 2021; Team, 2024; Wang et al., 2024b; Tian et al., 2024) and training with a next-token prediction objective. However, approaching the continuous distribution requires huge vocabulary sizes and a high utilization rate (Yu et al., 2023b; Weber et al., 2024), which is a complex objective.
@@
ACDiT is easy to implement, as simple as adding a tailored Skip-Causal Attention Mask to the current diffusion transformer (Peebles and Xie, 2022) during training. The inference process is formatted as an iteration between the conditional diffusion denoising process within a block, conditioned on the complete clean context, and autoregressive generation of a new block appended as the new context. In this way, KV-Cache can be used for faster inference. In general, ACDiT offers the following inherent advantages: (i) ACDiT simultaneously learns the causal interdependence across visual blocks with autoregressive modeling and the non-causal dependence within blocks with diffusion modeling. (ii) ACDiT is endowed with clean and continuous visual input, which eliminates the requirement for vector quantization and improves the transfer to visual understanding tasks after being trained with generative objectives. (iii) ACDiT makes full use of KV-Cache for flexible autoregressive generation of any length and can potentially benefit from other latest techniques in text for long video generation.
The effectiveness of ACDiT is primarily validated on visual generation tasks, including both image and video generation tasks, respectively. To further demonstrate its versatility, we also perform experiments on text generation. Experimental results demonstrate that ACDiT outperforms all autoregressive baselines of the same model scale, and achieves visual quality comparable to full-sequence diffusion models while exhibiting higher inference speed when extended to long sequences. It also achieves strong performance on text generation following the discrete diffusion formulation without any specific customization. We hope ACDiT presents a fresh perspective on the visual autoregressive model and paves the way for exploring unified multimodal and world models. The codes and models are available at https://github.com/thunlp/ACDiT.
@@ 2 Related Work
Diffusion Models. The field of image generation has witnessed remarkable advancements with the introduction of diffusion models (Ho et al., 2020; Song et al., 2020a; Dhariwal and Nichol, 2021; Nichol and Dhariwal, 2021). U-Net (Ronneberger et al., 2015) is the early mainstream choice of network architecture (Song et al., 2020b; Nichol and Dhariwal, 2021; Rombach et al., 2022; Podell et al., 2023). Following that, Transformer (Vaswani et al., 2017) is applied to diffusion models for image generation, with groundbreaking work such as DiT (Peebles and Xie, 2022) and U-ViT (Bao et al., 2023). A series work, including PixArt-$\alpha,\delta,\Sigma$(Chen et al., 2023; 2024c; 2024b), demonstrate the capability of DiT on text-to-image tasks. Several studies also apply DiT to video generation, such as Lumiere (Bar-Tal et al., 2024) and Movie Gen (Polyak et al., 2024).
@@
Autoregressive Generation on Discrete Tokens. The iGPT (Chen et al., 2020) first proposes autoregressively generating raw image pixels as a raster-scan sequence. Most visual autoregressive models (Ramesh et al., 2021; Wang et al., 2024b; Yu et al., 2022; Ding et al., 2021; 2022) follow VQGAN (Esser et al., 2021), which pioneers this direction by training an autoregressive transformer on discrete tokens produced by VQVAE (Van Den Oord et al., 2017). LlamaGen (Sun et al., 2024) enhances the image tokenizer and scales up the autoregressive transformers building on the latest Llama architecture (Touvron et al., 2023a). Inspired by RQ-Transformer (Lee et al., 2022), VAR (Tian et al., 2024) proposes the next-scale prediction and obtains good improvement. Some subsequent works (Tian et al., 2024; Li et al., 2024c; Yao et al., 2024) extend next-scale prediction to text-to-image generation.
@@
Autoregressive Generation on Continuous Features. Some recent works make some early exploration on achieving visual autoregressive generation without discrete tokens and combining the advantages of diffusion and autoregressive. Diffusion Forcing (Chen et al., 2024a) trains a causal autoregressive model to generate blocks without fully diffusing past ones and implements it on small RNN. MAR (Li et al., 2024a) proposes the diffusion loss to learn the autoregressive conditional distribution on the head of the main Transformer with a small MLP network. Transfusion (Zhou et al., 2024a) and Monoformer (Zhao et al., 2024a) conduct joint training with language modeling loss and diffusion loss in a single transformer, while within image generation, their design is identical to the standard diffusion model. Dart (Gu et al., 2024) proposes to autoregressively denoise the complete image within in a Transformer. Different from these works, ACDiT redefines the autoregressive unit and generates each block based on the clear past. Block Diffusion (Arriola et al., 2025) also explores an interpolation between autoregressive and diffusion modeling. Their approach focuses solely on text generation, building on discrete diffusion objectives proposed in prior work (Sahoo et al., 2024; Shi et al., 2024), and introduces variable-length generation and KV-caching within diffusion language models. In contrast, ACDiT is designed for visual generation, and is motivated by enabling autoregressive modeling over continuous visual data, yet still achieves better results on text generation using our simple and generally effective model architecture.
3 Prerequisite
3.1 Autoregressive Modeling
Autoregression asserts that the value at each timestep is contingent upon its preceding values. This principle is exemplified in autoregressive language models, which iteratively predict the probability distribution of subsequent tokens. Given a sequence of tokens ($x_{1},x_{2},\dots,x_{n}$), a salient characteristic of autoregression is that the prediction of $x_{i}$ is only dependent on its prefix ($x_{1},x_{2},\dots,x_{i-1}$). Upon determining $x_{i}$, it is concatenated with the preceding sequence, thereby forming the conditioning context ($x_{1},x_{2},\dots,x_{i}$) to predict $x_{i+1}$. Therefore, the likelihood of sequence can be factorized as:
$p(x_{1},x_{2},\dots,x_{N})=\prod_{i=1}^{N}p(x_{i}|x_{<i}).$ (1)
Thanks to the flexibility of self-attention in Transformers, autoregressive models can be effectively implemented by adding a causal attention mask in the Transformer attention block (Vaswani et al., 2017).
3.2 Diffusion
Diffusion models, in contrast, conceptualize a noise-infusion and denoising process, which is defined by gradually adding noise to the initial data $x_{0}\sim p(x)$ and training the model to learn the inverse mapping. Formally, the noised data $x^{(t)}$ at each step $t$ is sampled by $q(x^{(t)}|x^{(t-1)})=\mathcal{N}(x^{(t)};\sqrt{\alpha^{(t)}}x^{(0)},(1-\alpha^{(t)})\mathbf{I})$ [^1] For clarification, we use subscript t to denote timesteps in autoregressive models and superscript (t) to denote the timesteps in diffusion models., which is equivalent to add a Gaussian noise to the previous samples $x^{(t)}=\sqrt{\alpha^{(t)}}x^{(t-1)}+\sqrt{1-\alpha^{(t)}}\epsilon^{(t)},\epsilon^{(t)}\sim\mathcal{N}(0,\mathbf{I})$. $p_{\theta}$ is trained to learn the reverse process $p_{\theta}(x^{(t-1)}|x^{(t)})=\mathcal{N}(\mu_{\theta}(x^{(t)}),\beta^{(t)}\mathbf{I})$. With the reparameterization trick, the network $\mu_{\theta}(x^{(t)})$ can be reparameterized as noise prediction network $\epsilon_{\theta}(x^{(t)})$ and the training objective can be simple as:
[…]
During the inference phase, the denoising process is initialized with a random Gaussian noise sample $x^{(T)}$, followed by $T^{\prime}$ denoising steps, ultimately yielding a single deterministic sample $\tilde{n}^{(0)}$ from its underlying distribution. Typically, the denoising process in diffusion models operates “in-place”, meaning that each new denoising step directly replaces the previous step’s input. This differs from autoregressive modeling, where the value of a subsequent step is appended to the existing sequence.
4 ACDiT
4.1 Desiderata for Autoregressive Diffusion Model
Rather than modeling the full data distribution $p(x)$ directly, autoregressive diffusion aims to learn the conditional distributions for the continuous visual feature $p(x_{i}|x_{<i})$ through a diffusion process, which is typically represented as categorical distributions over discrete token vocabulary in language models. Combining the Eq.(1) and Eq.(2), the learning objective of autoregressive diffusion is:
[…]
where $\{x_{1},x_{2},...,x_{N}\}$ are $N$ autoregressive units. In the context of visual generation, each unit can flexibly correspond to the continuous visual information of a patch of pixels in an image or a set of frames in a video, depending on the granularity of the representation.
The key difference between Eq.(3) and Eq.(2) is that for each input noisy block, the noise prediction network $\epsilon(x_{i}^{(t)})$ predicts the noise conditioned on their previous clean block. To effectively learn this objective, we need a framework that integrates the strengths of both autoregressive and diffusion models. We identify three critical desiderata the framework should meet:
  1. 1.
  2. 2.
    Both the autoregressive modeling and the denoising process should optimally utilize the entire parameter space of the neural network. In an elegant fusion of autoregressive models and diffusion, neither component should be relegated to an auxiliary role. Instead, they should function as integral and complementary elements of the system.
  3. 3.
    The denoising process should directly attend comprehensively to the entire sequence of past sequences. Otherwise, the conditioning input to the denoising process would be required to fully encapsulate all prior information, imposing an unrealistic burden on the bottleneck context vector to losslessly compress the context feature into one vector. A holistic input of all past information in each denoise step ensures a more effective processing of temporal dependencies.
Based on these desiderata, we analyze representative autoregressive diffusion methods. Diffusion-Forcing (Chen et al., 2024a) proposes to use different-level random noise in different positions of a sequence. Thus, in the inference process, denoising subsequent positions from a clean past can be seen as a special case of denoising different levels of noise. Their method does not meet the first desideratum. In the training process, the future is not predicted from the precise representation of the past. In MAR (Li et al., 2024a), the diffusion process is trained based on the block in the last position, which does not satisfy the third desideratum. Moreover, the diffusion process sorely leverages the head part of the network, which conflicts with the second desideratum. To address suboptimal performance due to the causal direction based on this design, MAR incorporates bidirectional attention, which requires recomputing attention at each step and prevents the use of KV-Cache during inference. In Transfusion (Zhou et al., 2024a) and Monoformer (Zhao et al., 2024a), the autoregressive generation is restricted to the textual modality and is not applied within the visual modality. Image generation in these models follows the standard diffusion process conditioned on preceding text inputs, without incorporating autoregressive prediction over visual patches. Moreover, in visual understanding tasks or multi-image generation settings, subsequent text and images attend to the noisy version of the preceding image, which conflicts with the first desideratum. To mitigate this issue, these models reduce the noise schedule, i.e., halving the maximum number of diffusion steps in 20% of image-captioning pairs, which is not an optimal strategy. See the visual comparison of these models in Fig. 3(a).
@@ 4.1 Desiderata for Autoregressive Diffusion Model
Figure 2: (a): For each noised block $n_{i}$, it can only attend previous clean block $\{c_{0},c_{1},\dots,c_{i-1}\}$ and itself. Each clean block $c_{i}$ only attends to the previous clean blocks. (b): ACDiT can effectively utilize the KV-Cache during inference. (c): the 3D view of ACDiT, where $B,L,T$ denote the block size, number of blocks, and denoising time steps, respectively. The darker color denotes higher noise.